# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 5.1.2 Fine-tuning with Differential Privacy
Differential Privacy (DP) introduces various challenges, notably due to the addition of noise. One such issue, the increased memory consumption from noise addition, is addressed by Ghost clipping (Li et al., 2021). To further mitigate these challenges, Behnia et al. (2022) introduce a new framework designed to reduce the overall amount of noise required to maintain privacy guarantees. This framework utilizes the Edgeworth accountant (Wang et al., 2022), which provides a "finite sample guarantee" unlike the typically asymptotic nature of DP guarantees. This is particularly advantageous for fine-tuning an LLM where the number of tuning steps is relatively limited. By calculating the necessary noise levels based on privacy parameters and the number of compositions for a specific dataset, this new framework significantly reduces noise impact—by up to 5.6%. Additionally, it achieves a performance improvement of up to 1.1% across various natural language understanding (NLU) tasks compared to previous benchmarks. This makes the approach relevant as it not only enhances privacy but also improves efficiency in DP applications.
@@
Although pretraining on public data followed by fine-tuning on private data has long been considered the standard and safest method for private model training, recent evidence suggests that even this approach carries certain risks. Feng & Tramèr (2024) recently showed how privacy backdoors can be injected in pre-trained models making them susceptible to different kinds of attacks. This research explores how neurons can be subtly manipulated to retain their primary function while gaining malicious capabilities. By enabling neurons to hold a gradient for a single input and "deactivate" to prevent overwriting during subsequent training, the model becomes vulnerable to gradient inversion attacks, similar to those used in federated learning models (see 3.3 for details). While such attacks typically require white-box access, the researchers also demonstrate attacks on black-box models. Specifically, they enable membership inference attacks (MIAs, see 3.2) by modifying units in the model’s second-to-last layer, which are uniquely activated by specific training data points. By adjusting the model’s activation, they can determine whether a sample was part of the training set by comparing logits before and after fine-tuning. Although the study only tests on MLPs, ViT, and BERT, and it remains unclear if these vulnerabilities apply to LLMs, it highlights a significant risk in the public pre-training and private fine-tuning paradigm: if the pre-trained model or its provider is untrusted, loose privacy budgets commonly assumed in the literature are unsafe. Any efforts to establish tighter privacy guarantees must account for the model’s integrity.
@@ 5.1.3 Parameter-Efficient Fine-Tuning with Differential Privacy
Modern large language models (LLMs) often contain billions of parameters, making even full fine-tuning impractical due to high computational demands, particularly when paired with noise injection for Differential Privacy (DP). Instead, Parameter Efficient Fine-Tuning (PEFT) methods like adapter tuning (Houlsby et al., 2019), compacter tuning (Mahabadi et al., 2021), prompt tuning (Lester et al., 2021), prefix tuning (Li & Liang, 2021), and notably LoRA (Hu et al., 2021), have become more prevalent. Around the same time LoRA came out, also Reparametrized Gradient Perturbation (RGP) was introduced (Yu et al., 2021b), although with less success. RGP is a PEFT technique specific to the field of DP. It enhances efficiency by reconfiguring each weight matrix into smaller matrices, optimizing gradient storage and computation. These techniques, which involve adapting a subset of the model’s parameters or incorporating external parameters, facilitate faster training and enhanced modularity. They allow the model to manage various tasks without requiring complete re-training. Remarkably, even with fewer parameters being updated, these methods often achieve performance that is comparable to or even surpasses full fine-tuning, depending on the specific task and dataset (Sun et al., 2023; Suri et al., 2023). In this section, we survey methods that apply Differential Privacy with PEFT techniques in order to reduce the computational burden of fine-tuning while keeping privacy standards. A synthetic overview of these methods is available at 5.
@@
Yu et al. (2021a) applied all the PEFT techniques just mentioned with DP to RoBERTa and GPT2 without any particular challenge. As a matter of fact, they achieve performances nearly equivalent to their public counterparts, especially with RoBERTa where many PEFT techniques with DP notably outperformed full fine-tuning. The study also highlights the efficiency of DP when combined with LoRA in terms of speed and memory use, underscoring the higher viability of private learning in a PEFT framework. This is done by using DP-SGD with LoRA
@@
In a continuation of similar research, Li et al. (2021) conduct tests with both RoBERTa and GPT2, employing DP-Adam, which is similar to DP-SGD but based on Adam, for updates and moment accumulation) and the Ghost Clipping technique. This introduction reduces memory usage, making the difference in efficiency between private and non-private learning negligible (see §5 for more info). At the same time, also the performance drop is proven to be negligible, making DP-PEFT a viable option. GPT2 is evaluated in tasks like table-to-text and chit-chat generation, where the performance of full fine-tuning and PEFT proves to be closely matched, with full fine-tuning showing a slight edge both with and without Differential Privacy. This indicates that PEFT can match the efficacy of full fine-tuning even in a private setting. Additionally, the results challenge previous assumptions seen in the work of Yu et al. (2021a), where PEFT’s superior performance was attributed to lesser noise addition. The findings here suggest that higher noise levels in full fine-tuning might not detrimentally affect the privacy/utility tradeoff as previously thought. This observation prompts further inquiry into the mechanisms of private learning, offering a pathway to deeper insights into how different fine-tuning methods impact model performance under privacy constraints.
@@
While LoRA remains the predominant technique for private tuning, alternative methods also demonstrate effectiveness. Li et al. (2023d) introduce RAPT, a privatization framework designed for prompt-tuning and prefix-tuning on a model that does not necessarily need to be locally run. This is made specifically to make it possible to privately train a remote model through prompt tuning APIs, like Nvidia NeMo. This involves adding noise to the prompts before they are sent, addressing the issue that direct prompt tuning on privatized data can significantly impair downstream performances. To counteract this, they propose a novel privatized token reconstruction task, which is trained concurrently with the downstream task. The experiments utilize BERT and T5 models across various datasets such as SST (Socher et al., 2013), QQP (Lili Jiang, 2017), and TP-UK (Hovy et al., 2015), with adjustable privacy parameters. Privacy is assessed through the defensive capabilities against both an embedding inversion attack (see §3.3.1) and an attribute inference attack, where both prompt-tuning and prefix-tuning show enhanced robustness. Utility tests reveal that performances suffer without the use of the privatized token reconstruction, which significantly boosts performance and confirms its efficacy in helping LLMs develop better representations.
@@ 5.2 Inference with Differential Privacy
The first class of DP-inference models is based on In-context learning (ICL) (Brown et al., 2020a) (see §2). Indeed, ICL may be used to increase privacy in LLMs since private data can be given to LLMs only at inference time, and thus, this private data should not be included in generalistic LLMs and can stay under the control of the owners holding the data rights. For example, if a medical decision on a specific patient should be taken by considering all patients in a specific hospital, these patients can be given as part of the prompt in in-context learning (ICL) approach. However, nothing guarantees that a malicious user can try to extract this personal data by using specific prompts following the general ICL prompt (see §3.1.2). This is an active area of research that aims to embed Differential Privacy mechanisms in inference, performing what is called differentially private in-context learning. The work presented here all follow a similar high-level structure represented in Figure 7: the original dataset is used subsampled or partitioned in several subsets; the subsamples are then combined with the user prompt and fed to an LLM; all the generated outputs are aggregated in a private fashion, producing a single differentially private prediction.
@@
To address the issue rising with ICL, Wu et al. (2023a) introduce a novel approach, called Private in-context learning, to ensure privacy in ICL operations by partitioning the private dataset into M subsets, each containing N samples. These subsets are used as N-shot examples for generating M different ICL prompts, which in turn produce responses for the next token prediction or output classification. Results from each prompt are then privately aggregated using techniques tailored to the specific task: (1) Report-Noisy-Max with Gaussian noise for text classification that ensures differential privacy; (2) aggregation methods for text generation. Two aggregation methods are explored: Embedding Space Aggregation, where outputs are mapped to an embedding space to compute and noise the mean output, followed by querying the LLM multiple times to find the closest match with cosine similarity, and Keyword Space Aggregation, which involves noisy top-k counting of words across outputs using mechanisms like the Joint Exponential Mechanism (Gillenwater et al., 2022) or Propose-Test-Release (Zhu & Wang, 2022). Despite operating within a low privacy budget (e.g., $\epsilon=1$), Private in-context learning (ICL) achieves performance comparable to non-private methods across various tasks such as text classification, document question answering, and text summarization. This method not only demonstrates minimal performance drop with added privacy but also offers a less computationally intense alternative to fine-tuning with DP, providing a practical, lightweight solution for enforcing privacy in LLMs.
@@
Tang et al. (2024) added another security level on top of the method of Wu et al. (2023a): their method generates samples with DP before giving the input as in-context learning to LLMs. Then, the rest is the same, that is, possibly dividing the private dataset into M subsets containing the N samples. Clearly, Tang et al. (2024) show that the distortion of the input private data samples has little influence with respect to target applications. The effectiveness of this method is examined over benchmarks on standard Text Classification and Information Extraction datasets. Performance results are presented as the mean and standard deviation across five runs. The private models demonstrate performance on par with non-private models, indicating that it is feasible to achieve private learning without sacrificing accuracy. The study also evaluates the performance of Gaussian and Exponential mechanisms in the aggregation phase, with the Gaussian mechanism consistently showing slightly better results. This reinforces the viability of using synthetic data for privacy-preserving ICL applications.
@@
To improve the level of security in these ICL-based approaches, Duan et al. (2023a) propose to use a teacher-student approach where real data are given to the teacher model in a way similar to Wu et al. (2023a). The teacher model is then used to give synthetic examples for the ICL of the student model. In this way, the student model, which is exposed to final users, does not have access to real data. The method is called PromptPATE and follows the same general workflow of the standard Private Aggregation of Teacher Ensembles (PATE), presented by Papernot et al. (2018): first, an ensemble of teacher models is trained, then private knowledge transfer is performed, and lastly, a student model is trained. All this training has to be intended as ICL. In this case, instead of training the teacher models, the private dataset is used to generate a set of disjoint prompts for the LLM, so teacher models (or the "flock of stochastic parrots") are actually models performing ICL with disjoint parts of the original dataset. The private knowledge transfer is basically a majority voting from the teacher models made in order to choose the next token prediction. The noisy majority voting is done with the Confident GNMAX algorithm (Papernot et al., 2018) just like in the standard PATE. The chosen token is added to the public input to create a single, differentially private example that can replace private data for ICL on a student model so that the student model does not actually need to undergo real training but can take advantage of the elasticity of prompting. Similarly to Tang et al. (2024), this approach also involves generating synthetic samples for later use in ICL. This method, too, achieves performance comparable to non-private methods, even under stringent privacy conditions.
@@
Similarly, Tong et al. (2024) propose InferDPT, which includes a local perturbation algorithm called RANTEXT (discussed in §5). The perturbed text is sent to a remote black-box LLM, and a local extraction module — a smaller model — processes the perturbed output, reconstructing it to ensure coherence, consistency, and alignment with the original prompt. Users then combine the original prompt and the perturbed generation results to produce the final output. InferDPT, using the RANTEXT mechanism, achieves text generation quality comparable to non-private GPT-4, maintaining high utility despite the added privacy-preserving perturbations.
@@ 5.3 Federated Learning with Differential Privacy
Differential Privacy and federated learning are two complementary techniques in privacy-preserving machine learning that tackle different aspects of data security. Federated learning (§2.2) is a distributed computation method that enables learning on private data from various sites without transferring the data offsite. However, this alone is not sufficient for data security, as gradient inversion attacks can potentially recover the original training data by intercepting gradients sent from clients to the server (check discussion in §3.3.2). While federated learning helps mitigate memorization issues (Thakkar et al., 2020), real privacy guarantees require Differential Privacy. In federated learning, Differential Privacy can be implemented in different ways: Figure 8 illustrates its application either locally (local Differential Privacy) or globally (central Differential Privacy).
@@
A clarification is needed: central and local models of Differential Privacy (DP) are not exclusively related to federated learning (FL). In central Differential Privacy, perturbation is applied at a global level, typically during aggregation. In contrast, local Differential Privacy applies perturbation locally before the data leaves the site. While DP learning often aligns with FL and involves perturbing gradients, DP was initially developed for the statistical analysis of databases, and both central and local models exist independently of learning processes. §5.3.2 provides several examples of techniques that fall under the local model of DP, where text perturbation is applied to create a noisy private dataset that can be shared later, while §5.3.1 contains examples of techniques of central Differential Privacy, where the noise is applied at aggregation time.
5.3.1 Federated Learning with Central DP
McMahan et al. (2018) introduce DP-FedAvg, a variation of the FedAVG algorithm by McMahan et al. (2016), incorporating the moments accountant (Abadi et al., 2016) to calculate the privacy budget used during training. Typically, FedAVG performs part of the training locally on a subset of clients selected based on specific criteria. The clients then send their updated gradients to the server, which averages them, updates the global model, and sends the updated model back to the clients. In DP-FedAvg, gradients are clipped locally but not obfuscated. The central server adds noise after receiving and aggregating the gradients to prevent inversion, aligning the algorithm with central DP since local clipping alone does not meet the Local DP criteria. This approach was tested by training an LSTM for next token prediction on a large dataset of Reddit posts described by Al-Rfou et al. (2016), using various noise levels and clipping strategies. The model’s performance was evaluated using AccuracyTop1, the probability that the predicted word is correct. The study demonstrated that, with sufficient data, it is possible to train a model with privacy guarantees without significantly sacrificing performance, provided the noise added is kept within reasonable limits.
Currently, large language models (LLMs) represent the state of the art due to their superior power, though they are much larger than the small LSTMs seen in previous studies. Wang et al. (2023) focus on using LLMs to enhance the privacy/utility tradeoff of local LMs in differentially private federated learning. Similar to Yu et al. (2023), which samples pre-training data based on its similarity to fine-tuning data, they propose a distribution matching algorithm to sample public data close to the private data distribution, significantly improving sample efficiency. They utilize DP-FTRL (Kairouz et al., 2021a), an online optimization algorithm combining FTL (follow the leader) and regularization, adding noise during local updates. The local LM is either a single-layer LSTM (Hochreiter & Schmidhuber, 1997) or Transformer (Vaswani et al., 2023) due to computational limits on edge devices, while the LLM used is LaMDA (Thoppilan et al., 2022). Their tests yielded several notable results. First, using a public tokenizer with a public vocabulary instead of a private one helps prevent privacy leakage. Second, pre-training local models on a public dataset (C4 dataset) for over a week of single GPU time improved next token prediction accuracy (tested on the private StackOverflow dev set). They also built a distillation corpus by saving the top-k logits with k nonzero entries from the teacher LLM’s next token predictions as a silver label dataset. Pre-training on this distillation corpus showed accuracy comparable to pre-training with the original dataset, even with just 1% distillation coverage. Furthermore, their new distribution matching technique allows using only 0.08% of the pre-training dataset to achieve performance equal to 1% distillation coverage, greatly enhancing sample efficiency.
@@ 5.3.1 Federated Learning with Central DP
Xu et al. (2023b) present and analyze twenty Gboard LMs trained for Next Word Prediction using DP-FTRL, asserting that all future Gboard models will have DP guarantees. In their DP-FTRL implementation, model updates are clipped and noise is added during global model updates, ensuring Central Differential Privacy. They also use adaptive clipping (Andrew et al., 2022), which adjusts the clip norm each round by privately estimating the model’s norm at a targeted quantile. However, adaptive clipping showed no significant improvement and was less robust in experiments with large report goals, sometimes causing catastrophic failure or no progress in the first 1000 rounds. Despite this, adaptive clipping can reduce hyperparameter tuning when the privacy budget permits. Public pre-training on C4, as explained by Wang et al. (2023), reduced the rounds needed to reach target utility by  1000 under the same noise multiplier, enhancing privacy guarantees. They also combined DP-FTRL with Secure Aggregation (Bonawitz et al., 2017), a cryptographic protocol ensuring the central server only accesses aggregated updates, not individual client updates. While SecAgg aids DP, it introduces significant computational slowdown and requires different system configurations for large report goals. The techniques for training language models with differential privacy in Gboard have produced models with stronger privacy guarantees than some established standards, such as those used by the US Census Bureau. In Europe, these differentially private neural network models outperform older N-gram models in utility. For English in the US and Portuguese in Brazil, the new models are on par with or slightly better than their non-private counterparts.
5.3.2 Federated Learning with Local DP
Youn et al. (2023) introduce RQM (Randomized Quantization Mechanism), a novel approach to enforce DP in federated settings with improved communication efficiency, used alongside Distributed SGD on a CNN for image classification on EMNIST. RQM enhances DP-SGD by eliminating the need for explicit noise addition, instead performing RQM after gradient clipping. In RQM, gradient updates are encoded in quantization levels, which are then randomly subsampled. A randomized rounding procedure maps continuous gradient values to the nearest subsampled quantization level. RQM falls under Renyi Differential Privacy (Mironov, 2017), a relaxation of the original differential privacy definition, placing the method under Local DP since RQM is performed on the clients. Tests show that RQM outperforms the Poisson Binomial Mechanism (Chen et al., 2022) (a traditional noise injection method) and noise-free clipped SGD, achieving better accuracy with stronger privacy guarantees.

### v2

@@ 5.1.2 Fine-tuning with Differential Privacy
Differential Privacy (DP) introduces various challenges, notably due to the addition of noise. One such issue, the increased memory consumption from noise addition, is addressed by Ghost clipping (Li et al., 2022). To further mitigate these challenges, Behnia et al. (2022) introduce a new framework designed to reduce the overall amount of noise required to maintain privacy guarantees. This framework utilizes the Edgeworth accountant (Wang et al., 2023), which provides a "finite sample guarantee" unlike the typically asymptotic nature of DP guarantees. This is particularly advantageous for fine-tuning an LLM where the number of tuning steps is relatively limited. By calculating the necessary noise levels based on privacy parameters and the number of compositions for a specific dataset, this new framework significantly reduces noise impact—by up to 5.6%. Additionally, it achieves a performance improvement of up to 1.1% across various natural language understanding (NLU) tasks compared to previous benchmarks. This makes the approach relevant as it not only enhances privacy but also improves efficiency in DP applications.
@@
Although pre-training on public data followed by fine-tuning on private data has long been considered the standard and safest method for private model training, recent evidence suggests that even this approach carries certain risks. Feng & Tramèr (2025) recently showed how privacy backdoors can be injected in pre-trained models, making them susceptible to different kinds of attacks. This research explores how neurons can be subtly manipulated to retain their primary function while gaining malicious capabilities. By enabling neurons to hold a gradient for a single input and "deactivate" to prevent overwriting during subsequent training, the model becomes vulnerable to gradient inversion attacks, similar to those used in federated learning models (see 3.3 for details). While such attacks typically require white-box access, the researchers also demonstrate attacks on black-box models. Specifically, they enable membership inference attacks (MIAs, see 3.2) by modifying units in the model’s second-to-last layer, which are uniquely activated by specific training data points. By adjusting the model’s activation, they can determine whether a sample was part of the training set by comparing logits before and after fine-tuning. Although the study only tests on MLPs, ViT, and BERT, and it remains unclear if these vulnerabilities apply to LLMs, it highlights a significant risk in the public pre-training and private fine-tuning paradigm: if the pre-trained model or its provider is untrusted, loose privacy budgets commonly assumed in the literature are unsafe. Any efforts to establish tighter privacy guarantees must account for the model’s integrity.
@@ 5.1.3 Parameter-Efficient Fine-Tuning with Differential Privacy
Modern large language models (LLMs) often contain billions of parameters, making even full fine-tuning impractical due to high computational demands, particularly when paired with noise injection for Differential Privacy (DP). Instead, Parameter Efficient Fine-Tuning (PEFT) methods like adapter tuning (Houlsby et al., 2019), compacter tuning (Karimi Mahabadi et al., 2021), prompt tuning (Lester et al., 2021), prefix tuning (Li & Liang, 2021), and notably LoRA (Hu et al., 2022), have become more prevalent. Around the same time LoRA came out, also Reparametrized Gradient Perturbation (RGP) was introduced (Yu et al., 2021), although with less success. RGP is a PEFT technique specific to the field of DP. It enhances efficiency by reconfiguring each weight matrix into smaller matrices, optimizing gradient storage and computation. These techniques, which involve adapting a subset of the model’s parameters or incorporating external parameters, facilitate faster training and enhanced modularity. They allow the model to manage various tasks without requiring complete re-training. Remarkably, even with fewer parameters being updated, these methods often achieve performance that is comparable to or even surpasses full fine-tuning, depending on the specific task and dataset (Sun et al., 2023; Suri et al., 2023). Because of the huge number of parameters in LLMs, they seem to suit particularly well these methods, even though there is no theoretical limitation in applying them to other models. In this section, we survey methods that apply Differential Privacy with PEFT techniques in order to reduce the computational burden of fine-tuning while keeping privacy standards. A synthetic overview of these methods is available at 5.
@@
Yu et al. (2022) applied all the PEFT techniques just mentioned with DP to RoBERTa and GPT2 without any particular challenge. As a matter of fact, they achieve performances nearly equivalent to their public counterparts, especially with RoBERTa, where many PEFT techniques with DP notably outperformed full fine-tuning. The study also highlights the efficiency of DP when combined with LoRA in terms of speed and memory use, underscoring the higher viability of private learning in a PEFT framework. This is done by using DP-SGD with LoRA.
@@
In a continuation of similar research, Li et al. (2022) conduct tests with both RoBERTa and GPT2, employing DP-Adam, which is similar to DP-SGD but based on Adam, for updates and moment accumulation) and the Ghost Clipping technique. This introduction reduces memory usage, making the difference in efficiency between private and non-private learning negligible (see §5 for more info). At the same time, also the performance drop is proven to be negligible, making DP-PEFT a viable option. GPT2 is evaluated in tasks like table-to-text and chit-chat generation, where the performance of full fine-tuning and PEFT proves to be closely matched, with full fine-tuning showing a slight edge both with and without Differential Privacy. This indicates that PEFT can match the efficacy of full fine-tuning even in a private setting. Additionally, the results challenge previous assumptions seen in the work of Yu et al. (2022), where PEFT’s superior performance was attributed to lesser noise addition. The findings here suggest that higher noise levels in full fine-tuning might not detrimentally affect the privacy/utility tradeoff as previously thought. This observation prompts further inquiry into the mechanisms of private learning, offering a pathway to deeper insights into how different fine-tuning methods impact model performance under privacy constraints.
@@
While LoRA remains the predominant technique for private tuning of language models, alternative methods also demonstrate effectiveness. Li et al. (2023e) introduce RAPT, a privatization framework designed for prompt-tuning and prefix-tuning on a model that does not necessarily need to be locally run. This is made specifically to make it possible to privately train a remote language model through prompt tuning APIs, like Nvidia NeMo. This involves adding noise to the prompts before they are sent, addressing the issue that direct prompt tuning on privatized data can significantly impair downstream performances. To counteract this, they propose a novel privatized token reconstruction task, which is trained concurrently with the downstream task. The experiments utilize BERT and T5 models across various datasets such as SST (Socher et al., 2013), QQP (Lili Jiang, 2017), and TP-UK (Hovy et al., 2015), with adjustable privacy parameters. Privacy is assessed through the defensive capabilities against both an embedding inversion attack (see §3.3.1) and an attribute inference attack, where both prompt-tuning and prefix-tuning show enhanced robustness. Utility tests reveal that performances suffer without the use of the privatized token reconstruction, which significantly boosts performance and confirms its efficacy in helping LLMs develop better representations.
@@ 5.2 Inference with Differential Privacy
The methods presented here, be them in the first or second class, are all specifically devised for LLMs as they take advantage of prompting.
@@
The first class of DP-inference models is based on In-context learning (ICL) (Brown et al., 2020) (see §2.1). Indeed, ICL may be used to increase privacy in LLMs since private data can be given to LLMs only at inference time. Thus, this private data should not be included in generalistic LLMs and can stay under the control of the owners holding the data rights. For example, if a medical decision on a specific patient should be taken by considering all patients in a specific hospital, these patients can be given as part of the prompt in in-context learning (ICL) approach. However, nothing guarantees that a malicious user can try to extract this personal data by using specific prompts following the general ICL prompt (see §3.1.2). This is an active area of research that aims to embed Differential Privacy mechanisms in inference, performing what is called differentially private in-context learning. The work presented here all follow a similar high-level structure represented in Figure 7: the original dataset is used subsampled or partitioned in several subsets; the subsamples are then combined with the user prompt and fed to an LLM; all the generated outputs are aggregated in a private fashion, producing a single differentially private prediction.
@@
To address the issue rising with ICL, Wu et al. (2024b) introduce a novel approach, called Private in-context learning, to ensure privacy in ICL operations by partitioning the private dataset into M subsets, each containing N samples. These subsets are used as N-shot examples for generating M different ICL prompts, which in turn produce responses for the next token prediction or output classification. Results from each prompt are then privately aggregated using techniques tailored to the specific task: (1) Report-Noisy-Max with Gaussian noise for text classification that ensures differential privacy; (2) aggregation methods for text generation. Two aggregation methods are explored: Embedding Space Aggregation, where outputs are mapped to an embedding space to compute and noise the mean output, followed by querying the LLM multiple times to find the closest match with cosine similarity, and Keyword Space Aggregation, which involves noisy top-k counting of words across outputs using mechanisms like the Joint Exponential Mechanism (Gillenwater et al., 2022) or Propose-Test-Release (Zhu & Wang, 2022). Despite operating within a low privacy budget (e.g., $\epsilon=1$), Private in-context learning (ICL) achieves performance comparable to non-private methods across various tasks such as text classification, document question answering, and text summarization. This method not only demonstrates minimal performance drop with added privacy but also offers a less computationally intense alternative to fine-tuning with DP, providing a practical, lightweight solution for enforcing privacy in LLMs.
@@
Tang et al. (2024) added another security level on top of the method of Wu et al. (2024b): their method generates samples with DP before giving the input as in-context learning to LLMs. Then, the rest is the same, that is, possibly dividing the private dataset into M subsets containing the N samples. Clearly, Tang et al. (2024) show that the distortion of the input private data samples has little influence with respect to target applications. The effectiveness of this method is examined over benchmarks on standard Text Classification and Information Extraction datasets. Performance results are presented as the mean and standard deviation across five runs. The private models demonstrate performance on par with non-private models, indicating that it is feasible to achieve private learning without sacrificing accuracy. The study also evaluates the performance of Gaussian and Exponential mechanisms in the aggregation phase, with the Gaussian mechanism consistently showing slightly better results. This reinforces the viability of using synthetic data for privacy-preserving ICL applications.
@@
To improve the level of security in these ICL-based approaches, Duan et al. (2024a) propose to use a teacher-student approach where real data are given to the teacher model in a way similar to Wu et al. (2024b). The teacher model is then used to give synthetic examples for the ICL of the student model. In this way, the student model, which is exposed to final users, does not have access to real data. The method is called PromptPATE, and follows the same general workflow of the standard Private Aggregation of Teacher Ensembles (PATE), presented by Papernot et al. (2018): first, an ensemble of teacher models is trained, then private knowledge transfer is performed, and lastly, a student model is trained. All this training has to be intended as ICL. In this case, instead of training the teacher models, the private dataset is used to generate a set of disjoint prompts for the LLM, so teacher models (or the "flock of stochastic parrots") are actually models performing ICL with disjoint parts of the original dataset. The private knowledge transfer is basically a majority voting from the teacher models made in order to choose the next token prediction. The noisy majority voting is done with the Confident GNMAX algorithm (Papernot et al., 2018) just like in the standard PATE. The chosen token is added to the public input to create a single, differentially private example that can replace private data for ICL on a student model so that the student model does not actually need to undergo real training but can take advantage of the elasticity of prompting. Similarly to Tang et al. (2024), this approach also involves generating synthetic samples for later use in ICL. This method, too, achieves performance comparable to non-private methods, even under stringent privacy conditions.
@@
Similarly, Tong et al. (2024) propose InferDPT, which includes a local perturbation algorithm called RANTEXT (discussed in §4.1.2). The perturbed text is sent to a remote black-box LLM, and a local extraction module — a smaller model — processes the perturbed output, reconstructing it to ensure coherence, consistency, and alignment with the original prompt. Users then combine the original prompt and the perturbed generation results to produce the final output. InferDPT, using the RANTEXT mechanism, achieves text generation quality comparable to non-private GPT-4, maintaining high utility despite the added privacy-preserving perturbations.
@@ 5.3 Federated Learning with Differential Privacy
Differential Privacy and federated learning are two complementary techniques in privacy-preserving machine learning that tackle different aspects of data security. Federated learning (§2.3) is a distributed computation method that enables learning on private data from various sites without transferring the data offsite. However, this alone is not sufficient for data security, as gradient inversion attacks can potentially recover the original training data by intercepting gradients sent from clients to the server (check discussion in §3.3.2). While federated learning helps mitigate memorization issues (Thakkar et al., 2021), real privacy guarantees require Differential Privacy. In federated learning, Differential Privacy can be implemented in different ways: Figure 8 illustrates its application either locally (local Differential Privacy) or globally (central Differential Privacy).
@@
A clarification is needed: central and local models of Differential Privacy (DP) are not exclusively related to federated learning (FL). In central Differential Privacy, perturbation is applied at a global level, typically during aggregation. In contrast, local Differential Privacy applies perturbation locally before the data leaves the site. While DP learning often aligns with FL and involves perturbing gradients, DP was initially developed for the statistical analysis of databases, and both central and local models exist independently of learning processes. §C.1 provides several examples of techniques that fall under the local model of DP, where text perturbation is applied to create a noisy private dataset that can be shared later, while §C contains examples of techniques of central Differential Privacy, where the noise is applied at aggregation time.
Unfortunately, at the moment of writing, the intersection of differential privacy and federated learning did not produce much that is very relevant for LLMs, mainly due to the size of the models, which makes difficult parameter sharing and updating, and due to data dimensionality, which is an important constraint for text. There is still some interesting research we came across, even though not very relevant for LLMs, hence we leave the discussion for that in the appendix at §7.
@@
The most notable example of central differential privacy is Google’s implementation in Gboard. Xu et al. (2023b) discuss and analyze twenty Gboard language models trained for Next Word Prediction using DP-FTRL (refer to §7 for a thorough explanation), indicating that all future Gboard models will incorporate differential privacy guarantees. In their DP-FTRL setup, model updates undergo clipping, and noise is applied during global model updates, ensuring Central Differential Privacy. Adaptive clipping, which adjusts the clip norm each round by privately estimating the model’s norm at a targeted quantile as described by Andrew et al. (2021), was also utilized. However, adaptive clipping did not show a significant improvement and was less robust during experiments with large report goals, sometimes causing catastrophic failure or a lack of progress in the first 1000 rounds. Despite these issues, adaptive clipping can help reduce hyperparameter tuning when the privacy budget allows. Public pre-training on C4, as explained by Wang et al. (2024b), reduced the rounds needed to achieve target utility by about 1000 under the same noise multiplier, thereby enhancing privacy guarantees. Additionally, DP-FTRL was combined with Secure Aggregation, a cryptographic protocol ensuring the central server only accesses aggregated updates and not individual client updates, as described by Bonawitz et al. (2017). While Secure Aggregation supports DP, it introduces considerable computational slowdown and requires different system configurations for large report goals. The techniques for training language models with differential privacy in Gboard have resulted in models with stronger privacy guarantees than some established standards, such as those used by the US Census Bureau. In Europe, these differentially private neural network models outperform older N-gram models in terms of utility. In the US for English and in Brazil for Portuguese, the new models are comparable to or slightly better than their non-private versions.
For the local model of differential privacy, most works (as discussed in §7) use it for telemetry data collection, which means that they use very low dimensionality data. Data and model dimensionality are very problematic in this setting. Luckily, some recent developments can mitigate these problems. There is a very recent work (Liu et al., 2024b) that shows how this model can be applied to an LLM thanks to LoRA. In this work, the authors developed a federated version of LoRA with Differential Privacy, which allows the sharing of a much smaller number of parameters while keeping the privacy guarantee. The method makes use of LoRA to locally train the model, performs clipping and noising at a local level and then uploads the weight updates to the central server. They use this method to fine-tune Llama-7B and ChatGLM-6B for general and medical question answering, respectively, using SlimPajama and a medical dataset crawled from two online healthcare services platforms. The results for several $\varepsilon$ and $\delta$ show that, once again, there is a degradation in performance when using DP, but it is mostly negligible, especially considering the privacy gain. Unfortunately, in this paper, there is no concrete experiment to test to which extent this method prevents training data leakage.